OpenStudy (anonymous):

PLZ help!! i m hlf wy n stck! How many mL of 0.1 M HCl are required to react completely with 1 g mixture of Na2CO3 and NaHCO3 containing equimolar amounts of both?

OpenStudy (anonymous):

Let the amount of Na2CO3 in the mixture be x g. Then, the amount of NaHCO3 in the mixture is (1 − x) g. Molar mass of Na2CO3 = 2 × 23 + 1 × 12 + 3 × 16 = 106 g mol−1 Number of moles Na2CO3 Molar mass of NaHCO3 = 1 × 23 + 1 × 1 × 12 + 3 × 16 = 84 g mol−1 Number of moles of NaHCO3 According to the question, ⇒ 84x = 106 − 106x ⇒ 190x = 106 ⇒ x = 0.5579 Therefore, number of moles of Na2CO3 = 0.0053 mol And, number of moles of NaHCO3 = 0.0053 mol HCl reacts with Na2CO3 and NaHCO3 according to the following equation.

OpenStudy (anonymous):

now wht?????